15m : 40m = 35m : 65m
15m : 40m = 35m : 65m
15m : 40m = 35m : 65m

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

The given statement is False.
Consider the two ratios 15m : 40m = 35m : 65m

15/40 = 35/65
15/40 is further simplified by dividing both numerator and denominator by 5
we get, = 3/8
35/65 is further simplified by dividing both numerator and denominator by 5
we get, = 7/13
Hence, 3/8≠7/13
Therefore, 15/40≠35/65


Thanks